An aircraft shell is fired vertically upward with an initial velocity of 500 m/s. Find the ff:

a) maximum height it can reach

b) time taken to reach that height



1
Expert's answer
2020-11-05T10:43:50-0500

(a) Let’s take the upwards as the positive direction. Then, we can find the maximum height from the kinematic equation:


"v_f^2=v_i^2+2gh,"

here, "v_f=0\ ms^{-1}" is the final velocity of the shell at the maximum height, "v_i=500\ ms^{-1}" is the initial velocity of the shell, "g=-9.8\ ms^{-2}" is the acceleration due to gravity, "h" is the height.

Then, we get:


"h=-\dfrac{v_i^2}{2g}=-\dfrac{(500\ ms^{-1})^2}{2\cdot (-9.8\ ms^{-2})}=12.75\ km."

(b) We can find the time that the shell takes to reach the maximum height from the kinematic equation:


"v_f=v_i+gt,""t=\dfrac{v_f-v_i}{g}=\dfrac{0\ ms^{-1}-500\ ms^{-1}}{-9.8\ ms^{-2}}=51\ s."

Answer:

(a) "h=12.75\ km."

(b) "t=51\ s."